# Usman Ahmed,a Mikael P. Johansson,a,b Susi Lehtola,a and Dage Sundholm∗,a

### v1

@@ Abstract
Hydrogen bonding is an important non-covalent interaction that plays a major role in molecular self-organization and supramolecular structures. It can be described accurately with ab initio quantum chemical wave function methods, which become computationally expensive for large molecular assemblies. Density functional theory (DFT) offers a better balance between accuracy and computational cost, and can be routinely applied to large systems. A large number of density functional approximations (DFAs) has been developed, but their accuracy depend on the application, necessitating benchmark studies to guide their selection for use in applications. Some of us have recently determined highly accurate hydrogen bonding energies of 14 quadruply hydrogen-bonded dimers by extrapolating coupled-cluster energies to the complete basis set limit as well as extrapolating electron correlation contributions with a continued-fraction approach [U. Ahmed et al., Phys. Chem. Chem. Phys., 2024, 26, 24470]. In this work, we study the reproduction of these bonding energies at the DFT level using 152 DFAs. The top ten functionals are composed of eight variants of the Berkeley functionals both with and without dispersion corrections, and two Minnesota 2011 functionals augmented with a further dispersion correction. We find the B97M-V functional with the non-local correlation functional replaced by an empirical D3BJ dispersion correction to be the best functional, while changes to the dispersion part in other Berkeley functionals lead to poorer performance in our study.
@@ 1 Introduction
In this work, we study how well various DFAs reproduce the hydrogen bonding in 14 quadruply hydrogen-bonded dimers. Half of the molecules exhibit the DDAA-AADD and the other half exhibits the DADA-ADAD hydrogen bonding motif. We consider 152 DFAs including the Hartree–Fock (HF) method. The employed dispersion correction schemes include Grimme’s empirical dispersion corrections (DFT-D348, DFT-D449) and a non-local (NL) correlation functional (VV10).50
@@ 2 Computational methods
The benchmark considers the 14 molecules of the study of Ahmed et al.1 that determined hydrogen-bonding energies for TPSSh-D3/def2-TZVPP optimized geometries at the DLPNO-CCSD(T) level of theory extrapolated to the complete basis set limit, as well as towards the Schrödinger limit using the continued-fraction (cf) scheme of Goodson51. We denote the reference level of theory as CCSD(T)-cf. The reference energies are taken from Ahmed et al.1 and are included in the ESI†along with the employed geometries and visualizations thereof.
@@
The calculations were performed using the default (75,302) exchange-correlation grid in Psi4, which uses 75 radial grid points placed following Treutler and Ahlrichs55, and an angular Lebedev grid with 302 points.56 The VV10 non-local correlation functional50 was likewise evaluated on the corresponding (50,146) default grid. The resolution-of-the-identity approximation was used in all calculations with the universal Hartree–Fock auxiliary basis sets.57
The 75 base DFAs considered in this work, which have been listed in Table 1, comprise 14 semi-local generalized gradient approximation (GGA) functionals, 21 global hybrid (GH) GGA functionals, 13 range-separated (RS) GGA functionals, 13 semi-local meta-GGA (mGGA) functionals, 8 GH mGGA functionals, and 5 RS mGGA functionals, as well as HF (a GH functional without a semi-local exchange-correlation part). All of the density functionals were evaluated with Libxc,23 as usual in Psi4. Versions 6.0.0 and 6.2.2 of the library were employed. We note that some functionals such as B3LYP, HSE03, and HSE06 have various definitions in the literature,58 which is avoided through the use of the standard implementations in Libxc. Table 1 also shows the empirical (DFT-D348, DFT-D449) dispersion corrections or non-local correlation (VV1050) additionally considered with some of the functionals. Various combinations of these dispersion models (or the lack thereof) with the 75 base DFAs yield the 152 DFAs investigated in this study.
@@
B97-D3 GGA 59a 59 $\omega$B97 RS GGA 60
BLYP GGA 48a 49 61 62, 63, 64 $\omega$B97X RS GGA 49 60
BOP GGA 48a 65 $\omega$B97X-Db RS GGA 66
BP86 GGA 35a 67 62, 68 $\omega$B97X-V RS GGA 69c 70
CHACHIYO GGA 71, 72 $\omega$PBE RS GGA 48a 73, 74
KT1 GGA 75 CAM-B3LYP RS GGA 48a 76
KT2 GGA 75 CAMh-B3LYP RS GGA 77
KT3 GGA 78 CAM-QTP-00 RS GGA 79
N12 GGA 80a 81 CAM-QTP-01 RS GGA 82
PBE GGA 48a 49 67 83, 84 CAM-QTP-02 RS GGA 85
revPBE GGA 48a 49 61 86 HSE03 RS GGA 48a 49 87, 88
rPBE GGA 48a 49 89 HSE06 RS GGA 48a 49 87, 88, 90
PW91 GGA 91a 49 92, 93 N12-SX RS GGA 80a 94
XLYP GGA 48a 49 95 B97M-V mGGA 69c 96
HF GH 48a 49 M06-L mGGA 97
B1LYP GH GGA 48a 49 98 revM06-L mGGA 99
B3LYP GH GGA 48a 49 61 100 M11-L mGGA 80a 101
revB3LYP GH GGA 102 MN12-L mGGA 80a 103
B3PW91 GH GGA 48a 61 104 MN15-L mGGA 105
B3P86 GH GGA 35a 104 MVS mGGA 106, 107
B97 GH GGA 48c 108 SCAN mGGA 48a 49 109
B97-1 GH GGA 35a 110 rSCAN mGGA 111
B97-2 GH GGA 35a 112 r2SCAN mGGA 48a 113 114, 115
B97-3 GH GGA 116 TASK mGGA 117, 118
BHandH GH GGA 119 TPSS mGGA 48a 49 61 120, 121
BHandHLYP GH GGA 62, 119 revTPSS mGGA 48a 49 122 107, 123
BMK GH GGA 48a 124 M06 GH mGGA 31
HCTH120 GH GGA 48a 125 revM06 GH mGGA 126
HCTH407 GH GGA 48a 127 M06-2X GH mGGA 31
MPW1PW GH GGA 48a 49 128 M08-HX GH mGGA 129
O3LYP GH GGA 48a 49 130, 131 M08-SO GH mGGA 129
PBE0 GH GGA 48a 49 67 132, 133 MN15 GH mGGA 48a 134
revPBE0 GH GGA 48a 49 61 86, 83, 84 TPSSh GH mGGA 48a 49 27
QTP17 GH GGA 135 revTPSSh GH mGGA 48a 49 136
SOGGA11-X GH GGA 80a 137 $\omega$B97M-V RS mGGA 69c 138
X3LYP GH GGA 48a 49 95 M06-SX RS mGGA 139
M11 RS mGGA 48a 140
revM11 RS mGGA 141
MN12-SX RS mGGA 80a 94
@@
    Empirical dispersion with Becke–Johnson142 3-body damping.
@@
    Empirical dispersion with Becke–Johnson142 2-body damping.
@@ 3.1 Basis-set convergence
Fig.  1: Basis-set convergence plot of the hydrogen-bonding energies, relative to B97M-D3/def2-QZVPPD values. The figures show an energy window of ±$\pm 0.5$ kJ/mol as horizontal dashed lines.
3.2 The best functionals
The functionals were ranked based on their mean absolute errors (MAE). The ranking is shown in Tables 2 and 3 together with the corresponding MAEs and maximum absolute errors (MAX). The lowest MAE is obtained with the B97M-D3 functional, which is the B97M-V functional where the VV10 nonlocal correlation functional (indicated by -V in the functional’s name) has been replaced with an empirical D3 dispersion correction; the original B97M-V functional is ranked 4th. The second-lowest MAE is achieved with the $\omega$B97M-V functional, which is a RS GGA functional again with VV10 dispersion. Replacing VV10 with the D3 dispersion correction drops the ranking to the 7th place.
Table 2: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the 76 highest ranked functionals. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The functionals are ordered in increasing MAE.
@@ 3.2 The best functionals
Table 3: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the functionals ranked from 77 to 152. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The functionals are ordered in increasing MAE.
@@
The functionals ranked $5^{\mathrm{th}}$ to $9^{\mathrm{th}}$ are all Berkeley $\omega$B97 (RS GGA) functionals with different dispersion correction terms. $\omega$B97X-D3 is the $\omega$B97X-V70 functional with a D3 Becke–Johnson142, 48 dispersion correction instead of VV1069, and it should not to be confused with the $\omega$B97X-D3 functional of Lin et al.143 which was not studied in this work.
@@
Five of the six studied Berkeley functionals are ranked among the 10 best performing DFAs. mGGA and global hybrid (GH) GGA functionals also appear among the 20 most accurate functionals for hydrogen-bonding energies. The best GGA functionals are ranked 17th (XLYP-D3), 19th (rPBE-D4), 21st (BLYP-D3) and 22nd (XLYP-D4).
Fig.  2: Violin plots showing the error distributions in the hydrogen-bonding energies (in kJ/mol) for the investigated DFAs. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1
The error distributions for the hydrogen-bonding energies of the 36 top functionals are shown as violin plots in Fig. 2, while violin plots for the remaining functionals are included in the supporting information (SI). The average errors in the hydrogen-bonding energies for the top functionals are close to zero, except for M11-D3, whose distribution is almost completely on the negative side, implying that it systematically underestimates the hydrogen bonding energies. Similar plots are shown for the functionals ranked 13th–24th in Fig. 2, where the distributions are broader, and the average is larger than zero for many functionals, which implies that the hydrogen-bonding energies are overestimated. The error distributions for the functionals ranked between 25th and 36th shown in Fig. 2 are broader and more scattered than for the higher-ranked functionals. The violin plots for the rest of the functionals are given in the ESI†.
@@ 3.3 Meta-GGA functionals
The best performing mGGA functionals are B97M-D3 (1st), M11-L-D3 (3rd) and B97M-V (4th). SCAN and revTPSS-D3 are mGGA functionals that also perform well, holding rank 12 and 13, respectively. SCAN exhibits worse accuracy when paired with the D3 (ranked 75th) or D4 (ranked 62th) dispersion, suggesting overbinding. In contrast, the revTPSS functional improves its performance significantly by adding the empirical dispersion correction term to the functional, the rank increasing from 125th to 13th (D3) or 15th (D4), which suggests that the hydrogen bonds are too weak at the revTPSS level, which is rectified by the dispersion correction. The rest of the mGGA functionals among the first third part of the ranking list includes rSCAN (33rd), MVS (35th) and r2SCAN (37th). However, by considering the dispersion interaction using the D3 correction term, the performance of r2SCAN drops from 37th to 46th and to 44th with D4. Common mGGA functionals such as M06-L (53rd) and MN15-L (97th) are in the second third part of the ranking list, whereas M11-L without dispersion (104th), revM06-L (109th), MN12-L (120th), TPSS (126th), revTPSS (128th) and TASK (152nd) are in last third of the ranking list. The performance of the TPSS, revTPSS, MN12-L, and MN12-L functionals improves significantly when adding the empirical dispersion term. The accuracy is in a few cases improved by using D4 instead of D3, whereas using the VV10 dispersion correction leads to a lower accuracy than obtained with the D3 correction.
@@ 3.4 Range-separated hybrid functionals
The Berkeley range-separated functionals perform well. The $\omega$B97M-V is the best RS mGGA functional in the 2nd place and the Minnesota M11 functional combined with the D3 correlation term is ranked 10th. The best of the rest of the RS mGGA functionals is revM11, which is ranked 50th. Three of the five RS GGA functionals from Berkeley are ranked among the 10 best functionals, while adding a dispersion correction to $\omega$B97X leads to worse performace and 42nd rank. The $\omega$B97X-D functional performs poorly in our study (119th), but this is likely caused by lack of dispersion in our calculations with this functional; see Table 1.

### v2

@@ Abstract
Hydrogen bonding is an important non-covalent interaction that plays a major role in molecular self-organization and supramolecular structures. It can be described accurately with ab initio quantum chemical wave function methods, which become computationally expensive for large molecular assemblies. Density functional theory (DFT) offers a better balance between accuracy and computational cost, and can be routinely applied to large systems. A large number of density functional approximations (DFAs) has been developed, but their accuracy depend on the application, necessitating benchmark studies to guide their selection for use in applications. Some of us have recently determined highly accurate hydrogen bonding energies of 14 quadruply hydrogen-bonded dimers by extrapolating coupled-cluster energies to the complete basis set limit as well as extrapolating electron correlation contributions with a continued-fraction approach [U. Ahmed et al., Phys. Chem. Chem. Phys., 2024, 26, 24470]. In this work, we study the reproduction of these bonding energies at the DFT level using 152 DFAs. The top ten density functional approximations are composed of eight variants of the Berkeley functionals both with and without dispersion corrections, and two Minnesota 2011 functionals augmented with a further dispersion correction. We find the B97M-V functional with the non-local correlation functional replaced by an empirical D3BJ dispersion correction to be the best DFA, while changes to the dispersion part in other Berkeley functionals lead to poorer performance in our study.
@@ 1 Introduction
[…]
@@ 2 Computational methods
The benchmark considers the 14 dimers of the study of Ahmed et al.1 that determined hydrogen-bonding energies for TPSSh-D3/def2-TZVPP optimized geometries at the DLPNO-CCSD(T) level of theory extrapolated to the complete basis set limit, as well as towards the Schrödinger limit using the continued-fraction (cf) scheme of Goodson51. We denote the reference level of theory as CCSD(T)-cf. The reference energies and molecular structures of the dimers reported in the ESI† are taken from Ahmed et al.1. The molecular structures of the monomers with the DADA and DDAA motifs are shown in Figs. 1 and 2, respectively.
(a) DADA01
(b) DADA02
(c) DADA03
(d) DADA04
(e) DADA05
(f) DADA06
(g) DADA08
Fig.  1: The monomers of the quadruple hydrogen-bonded dimers with DADA-ADAD hydrogen-bonding pattern. The previously studied DADA071 is omitted because the CCSD(T)-cf reference data are not available. The figures are made with Jmol52.
(a) DDAA02
(b) DDAA03
(c) DDAA04
(d) DDAA05
(e) DDAA06
(f) DDAA07
(g) DDAA08
Fig.  2: The monomers of the quadruple hydrogen-bonded dimers with the DDAA-AADD hydrogen-bonding pattern. The previously studied DDAA011 is omitted because the CCSD(T)-cf reference data are not available. The figures are made with Jmol52.
@@
The hydrogen-bonding energies were corrected for basis set superposition errors (BSSE) calculated using the counterpoise correction method56, 57. The BSSE was obtained by comparing the energies of the two monomers of the dimer structure calculated using the dimer basis set to the monomer energies calculated using the monomer basis sets and the same molecular structure. The BSSE was added to the hydrogen-bonding energy calculated for fully optimized monomers using the basis set of the monomers.
The calculations were performed using the default (75,302) exchange-correlation grid in Psi4, which uses 75 radial grid points placed following Treutler and Ahlrichs58, and an angular Lebedev grid with 302 points.59 The VV10 non-local correlation functional50 was likewise evaluated on the corresponding (50,146) default grid. The resolution-of-the-identity approximation was used in all calculations with the universal Hartree–Fock auxiliary basis sets.60
The 75 base DFAs considered in this work, which have been listed in Table 1, comprise 14 semi-local generalized gradient approximation (GGA) functionals, 21 global hybrid (GH) GGA functionals, 13 range-separated (RS) GGA functionals, 13 semi-local meta-GGA (mGGA) functionals, 8 GH mGGA functionals, and 5 RS mGGA functionals, as well as HF (a GH functional without a semi-local exchange-correlation part). All of the density functionals were evaluated with Libxc,23 as usual in Psi4. Versions 6.0.0 and 6.2.2 of the library were employed. We note that some functionals such as B3LYP, HSE03, and HSE06 have various definitions in the literature,61 which is avoided through the use of the standard implementations in Libxc. Table 1 also shows the empirical (DFT-D348, DFT-D449) dispersion corrections or non-local correlation (VV1050) additionally considered with some of the functionals. Various combinations of these dispersion models (or the lack thereof) with the 75 base DFAs yield the 152 DFAs investigated in this study.
The study by Ahmed et al.1 showed that solvent effects increase the hydrogen-bonding energy of the tautomers DADA08 and DDAA06 by only 4-5%. We have not considered any solvent effects in this work because they are relatively small. Solvent effects are not expected to significantly affect the ranking of the DFAs.
@@
B97-D3 GGA 62a 62 $\omega$B97 RS GGA 63
BLYP GGA 48a 49 64 65, 66, 67 $\omega$B97X RS GGA 49 63
BOP GGA 48a 68 $\omega$B97X-Db RS GGA 69
BP86 GGA 35a 70 65, 71 $\omega$B97X-V RS GGA 72c 73
CHACHIYO GGA 74, 75 $\omega$PBE RS GGA 48a 76, 77
KT1 GGA 78 CAM-B3LYP RS GGA 48a 79
KT2 GGA 78 CAMh-B3LYP RS GGA 80
KT3 GGA 81 CAM-QTP-00 RS GGA 82
N12 GGA 83a 84 CAM-QTP-01 RS GGA 85
PBE GGA 48a 49 70 86, 87 CAM-QTP-02 RS GGA 88
revPBE GGA 48a 49 64 89 HSE03 RS GGA 48a 49 90, 91
rPBE GGA 48a 49 92 HSE06 RS GGA 48a 49 90, 91, 93
PW91 GGA 94a 49 95, 96 N12-SX RS GGA 83a 97
XLYP GGA 48a 49 98 B97M-V mGGA 72c 99
HF GH 48a 49 M06-L mGGA 100
B1LYP GH GGA 48a 49 101 revM06-L mGGA 102
B3LYP GH GGA 48a 49 64 103 M11-L mGGA 83a 104
revB3LYP GH GGA 105 MN12-L mGGA 83a 106
B3PW91 GH GGA 48a 64 107 MN15-L mGGA 108
B3P86 GH GGA 35a 107 MVS mGGA 109, 110
B97 GH GGA 48c 111 SCAN mGGA 48a 49 112
B97-1 GH GGA 35a 113 rSCAN mGGA 114
B97-2 GH GGA 35a 115 r2SCAN mGGA 48a 116 117, 118
B97-3 GH GGA 119 TASK mGGA 120, 121
BHandH GH GGA 122 TPSS mGGA 48a 49 64 123, 124
BHandHLYP GH GGA 65, 122 revTPSS mGGA 48a 49 125 110, 126
BMK GH GGA 48a 127 M06 GH mGGA 31
HCTH120 GH GGA 48a 128 revM06 GH mGGA 129
HCTH407 GH GGA 48a 130 M06-2X GH mGGA 31
MPW1PW GH GGA 48a 49 131 M08-HX GH mGGA 132
O3LYP GH GGA 48a 49 133, 134 M08-SO GH mGGA 132
PBE0 GH GGA 48a 49 70 135, 136 MN15 GH mGGA 48a 137
revPBE0 GH GGA 48a 49 64 89, 86, 87 TPSSh GH mGGA 48a 49 27
QTP17 GH GGA 138 revTPSSh GH mGGA 48a 49 139
SOGGA11-X GH GGA 83a 140 $\omega$B97M-V RS mGGA 72c 141
X3LYP GH GGA 48a 49 98 M06-SX RS mGGA 142
M11 RS mGGA 48a 143
revM11 RS mGGA 144
MN12-SX RS mGGA 83a 97
@@
    Empirical dispersion with Becke–Johnson145 3-body damping.
@@
    Empirical dispersion with Becke–Johnson145 2-body damping.
@@ 3.1 Basis-set convergence
Fig.  3: Basis-set convergence plot of the hydrogen-bonding energies, relative to B97M-D3/def2-QZVPPD values. The figures show an energy window of ±$\pm 0.5$ kJ/mol as horizontal dashed lines.
3.2 The best density functional approximations
The density functional approximations were ranked based on their mean absolute errors (MAE). The ranking is shown in Tables 2 and 3 together with the corresponding MAEs and maximum absolute errors (MAX). The lowest MAE is obtained with the B97M-D3 functional, which is the B97M-V functional where the VV10 nonlocal correlation functional (indicated by -V in the functional’s name) has been replaced with an empirical D3 dispersion correction; the original B97M-V functional is ranked 4th. The second-lowest MAE is achieved with the $\omega$B97M-V functional, which is a RS GGA functional again with VV10 dispersion. Replacing VV10 with the D3 dispersion correction drops the ranking to the 7th place.
Table 2: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the 76 highest ranked density functional approximations. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The DFAs are ordered in increasing MAE.
@@ 3.2 The best density functional approximations
Table 3: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the density functional approximations ranked from 77 to 152. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The DFAs are ordered in increasing MAE.
@@
The DFAs ranked $5^{\mathrm{th}}$ to $9^{\mathrm{th}}$ are all Berkeley $\omega$B97 (RS GGA) functionals with different dispersion correction terms. $\omega$B97X-D3 is the $\omega$B97X-V73 functional with a D3 Becke–Johnson145, 48 dispersion correction instead of VV1072, and it should not to be confused with the $\omega$B97X-D3 functional of Lin et al.146 which was not studied in this work.
@@
Five of the six studied Berkeley functionals are ranked among the 10 best performing DFAs. mGGA and global hybrid (GH) GGA functionals also appear among the 20 most accurate functionals for hydrogen-bonding energies. The best DFAs at the GGA level are ranked 17th (XLYP-D3), 19th (rPBE-D4), 21st (BLYP-D3) and 22nd (XLYP-D4).
Fig.  4: Violin plots showing the error distributions in the hydrogen-bonding energies (in kJ/mol) for the investigated DFAs. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1
The error distributions for the hydrogen-bonding energies of the 36 top DFAs are shown as violin plots in Fig. 4, while violin plots for the remaining DFAs are included in the ESI†. The average errors in the hydrogen-bonding energies for the top DFAs are close to zero, except for M11-D3, whose distribution is almost completely on the negative side, implying that it systematically underestimates the hydrogen bonding energies. Similar plots are shown for the DFAs ranked 13th–24th in Fig. 4, where the distributions are broader, and the average is larger than zero for many DFAs, which implies that the hydrogen-bonding energies are overestimated. The error distributions for the DFAs ranked between 25th and 36th shown in Fig. 4 are broader and more scattered than for the higher-ranked DFAs. The violin plots for the rest of the DFAs are given in the ESI†.
Hydrogen-bonding energies have been previously calculated in the basis-set limit at the second-order Møller-Plesset perturbation theory (MP2) level20. Since all deviations have the same sign, MP2 systematically overestimates the hydrogen-bonding energies, whereas the deviations of the hydrogen-bonding energies calculated using the top DFAs are both positive and negative. The hydrogen-bonding energies at the MP2 level correlate linearly with the reference energies. The angular coefficient of the linear fit of the MP2 energies is $1.022\pm 0.02$ with an offset of $-5.4\pm 3.3$ kJ/mol. The MAE and MAX values at the MP2 level are 8.3 kJ/mol and 14.6 kJ/mol, respectively, corresponding to the 55th place in the ranking list.
@@ 3.3 Meta-GGA functionals
The best performing DFAs at the mGGA level are B97M-D3 (1st), M11-L-D3 (3rd) and B97M-V (4th). SCAN and revTPSS-D3 are mGGA functionals that also perform well, holding rank 12 and 13, respectively. SCAN exhibits worse accuracy when paired with the D3 (ranked 75th) or D4 (ranked 62th) dispersion, suggesting overbinding. In contrast, the revTPSS functional improves its performance significantly by adding the empirical dispersion correction term to the functional, the rank increasing from 125th to 13th (D3) or 15th (D4), which suggests that the hydrogen bonds are too weak at the revTPSS level, which is rectified by the dispersion correction. The rest of the mGGA functionals among the first third part of the ranking list includes rSCAN (33rd), MVS (35th) and r2SCAN (37th). However, by considering the dispersion interaction using the D3 correction term, the performance of r2SCAN drops from 37th to 46th and to 44th with D4. Common mGGA functionals such as M06-L (53rd) and MN15-L (97th) are in the second third part of the ranking list, whereas M11-L without dispersion (104th), revM06-L (109th), MN12-L (120th), TPSS (126th), revTPSS (128th) and TASK (152nd) are in last third of the ranking list. The performance of the TPSS, revTPSS, MN12-L, and MN12-L functionals improves significantly when adding the empirical dispersion term. The accuracy is in a few cases improved by using D4 instead of D3, whereas using the VV10 dispersion correction leads to a lower accuracy than obtained with the D3 correction.
@@ 3.4 Range-separated hybrid functionals
The Berkeley range-separated functionals perform well. The $\omega$B97M-V is the best RS mGGA functional in the 2nd place and the Minnesota M11 functional combined with the D3 correlation term is ranked 10th. The best of the rest of the RS mGGA functionals is revM11, which is ranked 50th. Three of the five RS GGA functionals from Berkeley are ranked among the 10 best DFAs, while adding a dispersion correction to $\omega$B97X leads to worse performace and 42nd rank. The $\omega$B97X-D functional performs poorly in our study (119th), but this is likely caused by lack of dispersion in our calculations with this functional; see Table 1.